A truck proceeds to be able to pass another truck. Its speed increases from 50 to 100 km/h in 4 seconds. What is its average acceleration?



1
Expert's answer
2021-09-21T07:21:40-0400
"a=\dfrac{v-v_0}{t}=\dfrac{27.77\ \dfrac{m}{s}-13.88\ \dfrac{m}{s}}{4\ s}=3.47\ \dfrac{m}{s^2}."